OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose endpoints are given. P1 P2 M (3, 5) (-7, -5)

OpenStudy (anonymous):

formula is (x1+x2)/2 same for y

OpenStudy (anonymous):

(3+(-7))/2 is x (5-5)/2 is y

OpenStudy (anonymous):

-2,0

OpenStudy (anonymous):

thats the midpoint?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

:o really.... darn i got something els..

OpenStudy (anonymous):

\[M=\left( \frac{x_1+x_2}{2},\ \frac{y_1+y_2}{2}\right)\]

OpenStudy (anonymous):

:O oh noes :O

OpenStudy (anonymous):

timo can u check this one

OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose endpoints are given. P1 P2 M (10, 6) (-4, 8) is it 4,6?

OpenStudy (anonymous):

or is it-6

OpenStudy (anonymous):

3,7 :(

OpenStudy (anonymous):

WHATS!

OpenStudy (anonymous):

yeah 10-4 is 6 halfe that is 3 6+8 is 14 half that is 7 so 3,7

OpenStudy (anonymous):

WHATS!

OpenStudy (anonymous):

P1 P2 M (10, 6) (-4, 8) x1, y1 x2,y2 (x1+x2)/2, (y1+y2)/2 (10-4)/2=3, (6+8)/2=7 3 , 7